OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (solomonzelman):

so far you know that it is hoing to be y=-2x+b slope is -2, however "b" or the y-intercept is unknown. Plug in your point (5, -1) into y=-2x+b ---> (-1)=-2(5)+b -1=-10+b add 10 to both sides. -1+10=-10+b+10 9=b so knowing that b=9 and slope=-2, YOU GET----> y=-2x+9 Good?

OpenStudy (anonymous):

Wasn't one of the answers :o A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1) Those are the answers

OpenStudy (solomonzelman):

oh, in point slope from.... then A.

OpenStudy (solomonzelman):

y=-2x+9 add 1 to both sides y+1=-2x+10

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

Anytime!